# Context-Based Meta Reinforcement Learning for Robust and Adaptable Peg-in-Hole Assembly Tasks

### v3

@@ Abstract
Autonomous assembly is an essential capability for industrial and service robots, with Peg-in-Hole (PiH) insertion being one of the core tasks. However, PiH assembly in unknown environments is still challenging due to uncertainty in task parameters, such as the hole position and orientation, resulting from sensor noise. Although context-based meta reinforcement learning (RL) methods have been previously presented to adapt to unknown task parameters in PiH assembly tasks, the performance depends on a sample-inefficient procedure or human demonstrations. Thus, to enhance the applicability of meta RL in real-world PiH assembly tasks, we propose to train the agent to use information from the robot’s forward kinematics and an uncalibrated camera. Furthermore, we improve the performance by efficiently adapting the meta-trained agent to use data from force/torque sensor. Finally, we propose an adaptation procedure for out-of-distribution tasks whose parameters are different from the training tasks. Experiments on simulated and real robots prove that our modifications enhance the sample efficiency during meta training, real-world adaptation performance, and generalization of the context-based meta RL agent in PiH assembly tasks compared to previous approaches.
@@ I Introduction
PiH assembly has been studied in the literature as being the core of robotic assembly tasks [5, 6]. However, PiH assembly with uncertain parameters such as the hole position and orientation, resulting from sensor reading noise, remains challenging [1]. Meta reinforcement learning (RL) offers a promising paradigm by enabling robots to rapidly adapt to new insertion tasks and varying conditions using prior experience. Specifically, context-based meta RL offers a robust solution by learning to infer the unknown parameters from task data and adapting the policy accordingly [7, 8].
@@
In this paper, (i) we propose the modification of the context data to use forwards kinematics and easily measurable data from an uncalibrated vision sensor to enhance the real-world adaptation performance to address challenges arising from limited visual feedback, e.g., due to sensor occlusions, we develop a framework that enables the meta-trained agent to use alternative sensory modalities, such as force/torque feedback, (iii) we present an adaptation procedure to out-of-distribution (OOD) tasks, with parameters different from those of the training tasks, achieving generalization and solving one of the main limitations of context-based meta RL algorithms [7].
@@
Figure 2: The meta RL agent receives the hole position estimated from noisy external sensor, the current peg position (middle), the distance moved due to the last action calculated from forward kinematics, the detected hole and peg features in the 2D images captured before and after the action (left), and the force/torque sensor reading from the robot. The context encoder (right) uses these data to estimate the unknown information about the task, i.e., the actual hole position, and adapt the policy which produces the incremental robot motion.
@@ II Related Work
Previous methods have used vision and force/torque sensors to estimate the hole pose and perform the assembly [1]. Vision sensors are widely used in assembly tasks [10, 11], however, they require precise calibration. That is why, several methods have proposed an additional heuristic-based search step [12] to compensate for the remaining error [13, 14, 15]. Although previous works have achieved high accuracy in detecting hole features in 2D images, the 3D hole pose has high residual error due to noisy depth sensing which is further excacerbated by the texture of the hole surface [16, 17]. Hence, we propose to adapt the meta RL agent to rely on the detected hole features in the 2D image only without the need for depth information or accurate camera calibration. Force/torque sensors [18, 19, 20] have also been used to estimate the hole pose. However, they suffer from sim-to-real gap [21, 22], which limits their usage in learning-based methods. Therefore, we propose a sample-efficient method that adapts the meta RL agent, trained in simulation, to effectively utilize force/torque sensor data using a limited amount of real-world samples.
RL has been proposed to learn the robot actions directly from raw camera observations for assembly tasks [23, 24], however, it depends on a specific position and configurations for the camera and environment. Meta RL has been used to perform assembly tasks with uncertainty in the hole pose and different peg shapes. In [4], the meta RL agent is trained in simulation on a distribution of tasks with different but known hole positions and consequently known rewards. The agent is then tested in the real world on tasks with an unknown hole position. Due to the absence of the dense reward in the real world, the agent uses a sample inefficient procedure to adapt to new tasks. In [9], the meta RL agent uses human demonstrations to adapt to tasks with small uncertainty in the hole pose and different peg shapes. However, human demonstrations may not always be available. Therefore, we adapt the meta RL agent to use data that can be easily measured in the real world using onboard sensors.
Adapting the meta RL agent to use data from a different distribution, i.e., sensor, has also been proposed in [25]. A new context encoder is trained using supervised learning to produce the same output as the meta-trained encoder. However, the meta-trained encoder uses privileged information about the task and the new encoder uses a sequence of observations and actions as input and they produce a deterministic latent representation as output.
@@ III Problem Description
This work considers PiH assembly tasks with unknown hole position (see Fig. 2). The goal is to train an agent that quickly infers the accurate hole position and controls the motion of a robotic manipulator to insert the peg grasped by the end-effector into the hole. We assume that an external sensor such as camera provides the estimated hole position, but only with an error of ±$\pm\delta$ in the horizontal plane, similar to [4]. Furthermore, we assume that the manipulator is equipped with an eye-in-hand vision sensor and a force torque sensor at the wrist for context adaptation.
@@ IV Approach
To solve the Peg-in-Hole (PiH) assembly task with unknown hole position, we utilize the context-based meta RL algorithm PEARL [8], which is well-suited for this problem as it enables efficient task inference by leveraging context data collected during interactions. In our case, the agent has to infer the actual hole position from the context data and adjusts its policy accordingly. In this section, we first provide an overview of the PEARL algorithm and then introduce our proposed modifications to improve training and adaption efficiency, real-world applicability, and generalization to OOD tasks.
@@ IV-A PEARL Algorithm
PEARL [8] infers the current task using context data $c=(o,a,o^{\prime},r)$, where $o$ is the observation, $o^{\prime}$ the next observation, $a$ the action, and $r$ the reward. As shown in Fig. 3, the context data is fed to the context encoder to produce the mean and variance of a Gaussian distribution. After each collected trajectory, the produced distributions from all context data collected so far are multiplied together to produce one posterior Gaussian distribution over the latent space. The standard deviation represents the agent’s uncertainty about the task. The latent variables sampled from this distribution are fed to the policy to adapt its behavior to the current task. The context encoder in PEARL is trained to reconstruct the state action Q-value and the policy is trained to minimize the KL-divergence with the distribution over the Q-values. This implicitly means that the latent representation in PEARL reconstructs the policy behavior. As a result, representations close to each other in the latent space produce similar actions and policy behaviors and vice-versa.
@@ IV-A1 Observation, Reward, and Action Definitions
[…]
@@
where $\Delta x$, $\Delta y$, and $\Delta z$ denote the incremental movement in the $x$, $y$, and $z$ directions, respectively. The robot moves first in the horizontal direction then moves vertically similar to [4, 11].
@@ IV-A2 PEARL for PiH Tasks
Due to the noise in the estimated hole position, the actual hole can be located at any position within a range of ±$\pm\delta$ mm around the estimated position in the horizontal plane. As shown in Fig. 4, different possible actual hole positions have different rewards. Therefore, we consider each possible actual hole position around the estimated one as a different task similar to [4]. However, since the observation is based on the estimated hole position, it remains the same across all tasks. Meta RL can be applied to learn how to quickly adapt to tasks with the same observation and action spaces but with different rewards and dynamics [8, 26, 7]. The dynamics, which is the next observation given the current observation and action, are unique inside the hole for each task. This is because the peg will move beneath the hole surface at a given position only if the actual hole is located at this position. Any different actual hole position, i.e., task, will cause collision and prevent the downwards motion of the peg.
@@ IV-B Context Data Modification
Since the reward cannot be measured in the real-world due to the unknown actual hole position, we propose to modify the context data and use data that can be easily measured in the real world using the robot’s forward kinematics and uncalibrated camera. Additionally, using the reward in the context data for PiH assembly tasks does not provide clear information about the task as explained next.
@@
(b) Posterior Gaussian distribution.
Figure 5: An example of tasks with the same reward and original context data [4] and the corresponding posterior Gaussian distribution. Tasks with the same original context data require opposite actions resulting in high variance posterior Gaussian distribution.
IV-B1 Effect of the Context Data on Task Inference
@@
(b) Posterior Gaussian distribution.
@@ IV-B2 Proposed Modified Context Data
We therefore propose to replace the reward in the context data with the motion towards the actual hole $m$ based on the previous action. The motion $m$ is the summation of the distances moved by the peg towards the actual hole position in the three Cartesian coordinates, and the angle rotated around the vertical axis towards the actual hole in case of additional uncertainty in the hole orientation. $m$ is defined as follows
$m=\pm d_{x}\pm d_{y}\pm d_{z}\pm a_{z},$ (3)
where $d_{x}$, $d_{y}$, and $d_{z}$ are the distances moved by the peg in X, Y, and Z directions, respectively, and $a_{z}$ is the angle rotated by the peg around the vertical axis. Each distance is positive if it decreases the distance between the peg and the actual hole and vice-versa. During training in simulation, the actual hole position is known and $m$ can be accurately calculated. During execution in the real world, the distance moved in each direction is calculated from the robot forward kinematics. The sign of each distance is positive if it decreases the distance, in terms of pixels, between the detected hole and peg features in the 2D images captured before and after the motion in that direction, as shown in Fig. 2, and negative otherwise. The sign of $d_{z}$ is determined based on the known vertical position of the hole. Note that this procedure does not require camera calibration and it is agnostic to the position of the camera as long as the peg and hole features can be detected before and after the motion in each direction.
IV-B3 Effect of the Modified Context on Task Inference
In addition to being easily measurable, using the motion $m$ in the context data instead of the reward results in that tasks with the same modified context data $c=(o,a,o^{\prime},m)$ are closer to each other in the latent space. For example, in Fig. 6(a), tasks A and B have the same modified context data. Since they require slightly different actions, their latent representations are close to each other in the latent space resulting in a narrower Gaussian distribution by the context encoder (see Fig. 6(b)), and the agent becomes more certain about the best behavior. In case of additional uncertainty in the hole orientation, tasks with the same modified context, as tasks C and D, may require opposite actions in one dimension, however, it remains more informative than the original context data with opposite action in three dimensions.
@@ IV-C Adaptation to Different Sensor Data
To enhance the agent’s ability to perform PiH assembly tasks where the peg occludes the hole or camera is not available, We adapt the meta-trained agent to use data from a force/torque sensor to perform the PiH assembly task by training an alternative context encoder.
@@ IV-C1 Training Data Collection
To collect the training data, we apply the meta-trained agent on a number of real-world tasks with different hole positions. Data collection is done using latent variables sampled from the latent space of the meta training tasks to ensure safe and diverse actions by the policy. The collected actual context data $(o,a,o^{\prime},m)$ and the alternative context data $(o,a,o^{\prime},f)$, where $f$ is the force/torque sensor readings, from each task are stored in separate buffers.
@@ IV-C2 Training Procedure
(d) Actual hole positions of the training and test tasks.
@@
Using the posterior Gaussian distributions for the loss calculation, rather than the output of each individual sample, enables the new encoder to assign different variance for each data sample in the batch. For example, non-informative samples with no contact or no overlapping between the peg and the hole should be assigned high variance by the encoder so that they barely affect the posterior distribution.
@@ IV-D Out-of-Distribution (OOD) Adaptation
[…]
$L_{\text{OOD}}=-\alpha_{1}\sum_{n=1}^{N}(prob(z_{n})\cdot m_{n})+\alpha_{2}\frac{1}{\sigma^{2}}+\alpha_{3}(\mu-\overline{\mu_{i}})^{2},$ (4)
where $N$ is the number of data sampled from the current posterior distribution, $z_{n}$ is the sampled latent variable at step $n$ and $m_{n}$ is the resulted motion towards the actual hole at the same step, $\mu$ and $\sigma^{2}$ are the mean and variance of the posterior Gaussian distribution produced by the context encoder using the recently collected N samples, $\overline{\mu_{i}}$ is the mean of the current posterior Gaussian distribution before applying the optimization step and the overline means no gradient, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are tuning weights. The first term in Eq. 4 increases the probability of latent variables with high motion towards the actual hole and vice versa, using this term alone may severely decrease the variance which negatively affects the exploration in the subsequent steps. Therefore, we add the second term to penalize the small variance. The third term prevents huge changes in the posterior distribution between successive steps to ensure safety and prevent unpredictable behavior.

### v4

@@ Abstract
Autonomous assembly is an essential capability for industrial and service robots, with Peg-in-Hole (PiH) insertion being one of the core tasks. However, PiH assembly in unknown environments is still challenging due to uncertainty in task parameters, such as the hole position and orientation, resulting from sensor noise. Although context-based meta reinforcement learning (RL) methods have been previously presented to adapt to unknown task parameters in PiH assembly tasks, the performance depends on a sample-inefficient procedure or human demonstrations. Thus, to enhance the applicability of meta RL in real-world PiH assembly tasks, we propose to train the agent to use information from the robot’s forward kinematics and an uncalibrated camera. Furthermore, we improve the applicability by efficiently adapting the meta-trained agent to use data from force/torque sensor. Finally, we propose an adaptation procedure for out-of-distribution tasks whose parameters are different from the training tasks. Experiments on simulated and real robots prove that our modifications enhance the sample efficiency during meta training, real-world adaptation performance, and generalization of the context-based meta RL agent in PiH assembly tasks compared to previous approaches.
@@ I Introduction
PiH assembly has been studied in the literature as being the core of robotic assembly tasks [5, 6]. However, PiH assembly with uncertain parameters such as the hole position and orientation, resulting from sensor noise, remains challenging [1]. Meta reinforcement learning (RL) offers a promising paradigm by enabling robots to rapidly adapt to new insertion tasks and varying conditions using prior experience. Specifically, context-based meta RL offers a robust solution by learning to infer the unknown parameters from task data and adapting the policy accordingly [7, 8].
@@
In this paper, (i) we propose the modification of the context data to use forward kinematics and easily measurable data from an uncalibrated vision sensor to enhance the real-world adaptation performance to address challenges arising from limited visual feedback, e.g., due to sensor occlusions, (ii) we develop a framework that enables the meta-trained agent to use alternative sensory modalities, such as force/torque feedback, (iii) we present an adaptation procedure to out-of-distribution (OOD) tasks, with parameters different from those of the training tasks, achieving generalization and solving one of the main limitations of context-based meta RL algorithms [7].
@@
Figure 2: The meta RL agent receives the hole position estimated from noisy external sensor, the current peg position (middle), the distance moved due to the last action calculated from forward kinematics, the detected hole and peg features in the 2D images captured before and after the action (left), and the force/torque sensor reading from the robot. The context encoder (right) uses these data to estimate the unknown information about the task, i.e., the actual hole position, and adapts the policy which produces the incremental robot motion.
@@ II Related Work
Vision [10, 11] and force/torque sensors have been used to estimate the hole pose and perform the assembly [1]. Vision sensors require precise calibration and an additional heuristic-based search step [12] to compensate for residual errors [13, 14, 15]. Although earlier works have achieved high accuracy in detecting hole features in 2D images, estimating the 3D hole pose suffers from high residual errors due to noisy depth sensing [16, 17]. Hence, we adapt the meta RL agent to solely rely on the hole features in the 2D image without requiring depth information or accurate camera calibration. Unlike traditional visual servoing [18], which requires approximate or estimated depth values and intrinsic parameters in addition to image features to control robot motion, our agent relies solely on image features. Furthermore, it leverages features detected over all previous time steps to control robot motion, which enhances robustness to instantaneous perturbations, such as occlusions occurring at certain times, compared to visual servoing that depends only on the current image features to decide the robot motion.
Force/torque sensors [19, 20, 21] have also been used to estimate the hole pose. However, they suffer from the sim-to-real gap [22, 23], which limits their usage in learning-based agents trained in simulation. Therefore, we propose a sample-efficient method that adapts the meta RL agent, trained in simulation, to effectively utilize force/torque sensor data using a limited amount of real-world samples.
RL has been proposed to learn robot actions directly from raw camera observations for assembly tasks [24, 25], however, its generalization to changes in camera position and task configurations, such as peg shape and color, is limited. That is why, we use the detected features rather than raw pixels to train the agent to ensure its generalization to different task configurations. Meta RL has been used to perform assembly tasks with uncertainty in the hole pose and different peg shapes. In [4], the meta RL agent is trained in simulation on a distribution of tasks with different but known hole positions and consequently known rewards. The agent is then tested in the real world on tasks with an unknown hole position. Due to the absence of the dense reward in the real world, the agent uses a sample inefficient procedure to adapt to new tasks. In [9], the meta RL agent uses human demonstrations to adapt to tasks with small uncertainty in the hole pose and different peg shapes. However, human demonstrations may not always be available. Therefore, we adapt the meta RL agent to use data that can be easily measured in the real world using onboard sensors.
Adapting the meta RL agent to use data from a different distribution, i.e., a different sensor, has also been proposed in [26]. A new context encoder is trained using supervised learning to produce the same output as the meta-trained encoder. However, the meta-trained encoder uses privileged information about the task and the new encoder uses a sequence of observations and actions as input and they produce a deterministic latent representation as output.
@@ III Problem Description
This work considers PiH assembly tasks with unknown hole position (see Fig. 2). The goal is to train an agent that quickly infers the accurate hole position and controls the motion of a robotic manipulator to insert the peg grasped by the end-effector into the hole. We assume that an external sensor such as a camera provides the estimated hole position, but only with an error of $\pm\delta$ in the horizontal plane, similar to [4]. Furthermore, we assume that the manipulator is equipped with an eye-in-hand vision sensor and a force torque sensor at the wrist for context adaptation.
@@ IV Approach
To solve the Peg-in-Hole assembly task with unknown hole position, we utilize the context-based meta RL algorithm PEARL (Probabilistic Embeddings for Actor-critic RL) [8], which is well-suited for this problem as it enables efficient task inference by leveraging context data collected during interactions. In our case, the agent has to infer the actual hole position from the context data and adjust its policy accordingly. In this section, we first provide an overview of the PEARL algorithm and then introduce our proposed modifications to improve training and adaption efficiency, real-world applicability, and generalization to OOD tasks.
@@ IV-A PEARL Algorithm
PEARL [8] infers the current task using context data $c=(o,a,o^{\prime},r)$, where $o$ is the observation, $o^{\prime}$ the next observation, $a$ the action, and $r$ the reward. As shown in Fig. 3, the context data is fed to the context encoder to produce the mean and variance of a Gaussian distribution. After each collected trajectory, the produced distributions from all context data collected so far are multiplied together to produce one posterior Gaussian distribution over the latent space. The standard deviation represents the agent’s uncertainty about the task. The latent variables sampled from this distribution are fed to the policy to adapt its behavior to the current task. The context encoder is trained to reconstruct the state action Q-value and the policy is trained to minimize the KL-divergence with the distribution over the Q-values. Since the context encoder reconstructs the Q-value that is used to train the policy, the context encoder is considered to implicitly reconstruct the policy’s behavior. As a result, representations close to each other in the latent space produce similar actions and policy’s behaviors and vice-versa.
@@ IV-A1 Observation, Reward, and Action Definitions
The action $\mathbf{a}$ is defined as an incremental movement of the peg up to $2\text{\,}\mathrm{m}\mathrm{m}$ in the three Cartesian coordinates
@@
where $\Delta x$, $\Delta y$, and $\Delta z$ denote the incremental movement in the $x$, $y$, and $z$ directions, respectively. The robot moves first in the horizontal direction and then it moves vertically, similar to [4, 11], to avoid friction between the peg and the hole surface, which eliminates the need for force control.
@@ IV-A2 PEARL for PiH Tasks
Due to the noise in the estimated hole position, the actual hole can be located at any position within a range of $\pm\delta$ around the estimated position in the horizontal plane. As shown in Fig. 4, different possible actual hole positions have different rewards. Therefore, we consider each possible actual hole position around the estimated one as a different task similar to [4]. However, since the observation is based on the estimated hole position, it remains the same across all tasks. Meta RL can be applied to learn how to quickly adapt to tasks with the same observation and action spaces but with different rewards and dynamics [8, 27, 7]. The dynamics, which is the next observation given the current observation and action, are unique inside the hole for each task. This is because the peg will move beneath the hole surface at a given position only if the actual hole is located at this position. Any different actual hole position, i.e., different task, will cause collisions and prevent the downwards motion of the peg.
@@ IV-B Context Data Modification
Since the reward cannot be measured in the real-world due to the unknown actual hole position, we propose to modify the context data and use data that can be easily measured in the real world using the robot’s forward kinematics and an uncalibrated camera. Additionally, using the reward in the context data for PiH assembly tasks does not provide clear information about the task as explained next.
@@
(b) Posterior Gaussian distribution inferred by the context encoder.
Figure 5: An example of tasks with the same reward and original context data [4] and the corresponding posterior Gaussian distribution. Tasks with the same original context data require opposite actions resulting in a high variance posterior Gaussian distribution.
IV-B1 Effect of the Context Data on the Task Inference
@@
(b) Posterior Gaussian distribution inferred by the context encoder.
@@ IV-B2 Proposed Modified Context Data
We therefore propose to replace the reward in the context data with the motion towards the actual hole $m$ based on the previous action. The motion $m$ is the summation of the distances moved by the peg towards the actual hole position in the three Cartesian coordinates, and the angle rotated around the vertical axis towards the actual hole in case of additional uncertainty in the hole orientation. $m$ is defined as follows:
$m=\pm d_{x}\pm d_{y}\pm d_{z}\pm\theta_{z},$ (3)
where $d_{x}$, $d_{y}$, and $d_{z}$ are the distances moved by the peg in X, Y, and Z directions, respectively, and $\theta_{z}$ is the angle rotated by the peg around the vertical axis. Each distance is positive if it decreases the distance between the peg and the actual hole and vice-versa. During training in simulation, the actual hole position is known and $m$ can be accurately calculated. During execution in the real world, the distance moved in each direction is calculated from the robot forward kinematics. The sign of each distance is positive if it decreases the distance, in terms of pixels, between the detected hole and peg features in the 2D images captured before and after the motion in that direction, as shown in Fig. 2, and negative otherwise. The sign of $d_{z}$ is determined based on the known vertical position of the hole. Note that this procedure does not require camera calibration and it is agnostic to the position of the camera as long as the peg and the hole features can be detected before and after the motion.
IV-B3 Effect of the Modified Context on the Task Inference
In addition to being easily measurable, using the motion $m$ in the context data instead of the reward results in that tasks with the same modified context data $c=(o,a,o^{\prime},m)$ are closer to each other in the latent space. For example, in Fig. 6(a), tasks A and B have the same modified context data. Since they require slightly different actions, their latent representations are close to each other in the latent space resulting in a narrower Gaussian distribution by the context encoder (see Fig. 6(b)), and the agent becomes more certain about the suitable behavior. In case of additional uncertainty in the hole orientation, tasks with the same modified context, as tasks C and D, may require opposite actions in one dimension, however, it remains more informative than the original context data with opposite action in three dimensions.
@@ IV-C Adaptation to Different Sensor Data
To enhance the agent’s ability to perform PiH assembly tasks where the peg occludes the hole or the camera is not available, we adapt the meta-trained agent to use data from a force/torque sensor to perform the PiH assembly task by training an alternative context encoder.
@@ IV-C1 Training Data Collection
To collect the training data, we apply the meta-trained agent on a number of real-world tasks with different hole positions. We use latent variables sampled from the latent space of the meta training tasks to ensure safe and diverse actions by the policy. The collected actual context data $(o,a,o^{\prime},m)$ and the alternative context data $(o,a,o^{\prime},f)$, where $f$ is the force/torque sensor readings, from each task are stored in separate buffers.
@@ IV-C2 Training Procedure
(d) Actual hole position of the tasks relative to the estimated position.
@@
Using the posterior Gaussian distributions for the loss calculation, rather than the output of each individual sample, enables the new encoder to assign a different variance to each data sample in the batch. For example, non-informative samples with no contact or no overlapping between the peg and the hole should be assigned a high variance by the encoder so that they barely affect the posterior distribution.
@@ IV-D Out-of-Distribution (OOD) Adaptation
Similar to [4], we consider an error of $\pm$$4\text{\,}\mathrm{m}\mathrm{m}$ in the actual hole position. Thus, tasks with more error in the actual hole position will be considered as OOD tasks. We adapt to OOD tasks by first sampling exploration data using the normal prior distribution over the latent space and inferring the posterior distribution. This step leverages the agent’s previous experience to achieve the best possible initial behavior in OOD tasks. Then, we apply an iterative process of collecting exploration data using latent variables sampled from the current posterior Gaussian distribution and optimizing the context encoder network to push the posterior Gaussian distribution towards latent representations with high values of motion towards the actual hole $m$. The context encoder is optimized to minimize the following loss:
$L_{\text{OOD}}=-\alpha_{1}\sum_{n=1}^{N}(prob(z_{n})\cdot m_{n})+\alpha_{2}\frac{1}{\sigma^{2}}+\alpha_{3}(\mu-\text{sg}(\mu_{i}))^{2},$ (4)
where $N$ is the number of data sampled from the current posterior distribution, $z_{n}$ is the sampled latent variable at step $n$ and $m_{n}$ is the resulted motion towards the actual hole at the same step, $\mu$ and $\sigma^{2}$ are the mean and variance of the posterior Gaussian distribution produced by the context encoder during the current optimization step using the recently collected N samples, $\mu_{i}$ is the mean of the posterior Gaussian distribution before applying the optimization step, i.e., the posterior distribution resulting from the previous optimization step, and sg denotes to stop gradient, and $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are tuning weights. The first term in Eq. 4 increases the probability of latent variables with high motion towards the actual hole and vice versa. Using this term alone may severely decrease the variance, in an attempt to increase the probability, which negatively affects the exploration in the subsequent steps. Therefore, we add the second term to penalize the small variance. The third term prevents large changes in the posterior distribution between successive optimization steps to ensure safety and stability and prevent unpredictable behavior.